### v1

@@ 1. Introduction
In recent years, the integration of large language models (LLMs) into conversational robots—embodied media—has enabled more natural human-robot conversations. However, while LLMs can mitigate certain speech recognition errors that often cause traditional conversational robots to fail, these advanced systems remain vulnerable to issues such as interrupting users and misinterpreting user intent (Mahmood et al., 2025b, 2024; Cao et al., 2025). Such failures can lead to conversational breakdowns, disrupt the interaction flow, and negatively affect the user’s perception of the robot. As LLM-powered conversational robots are being deployed in real-world scenarios, the need to understand and mitigate their failures is a pressing issue. Previous works have discovered some strategies to help robots rebuild user trust and recover from errors (Karli et al., 2023; Esterwood and Robert, 2021); however, the application of these strategies hinges on the robot’s ability to detect its own errors effectively (Stiber et al., 2023). Status quo robot error detection methods often either depend on task-specific or domain-specific information—which can be hard to generalize across people, tasks, and errors—or rely on users to detect and report errors—which can be delayed. Recent works showed the possibility of detecting robot error through leveraging social signals exhibited by user reactions and responses to robot errors in multimedia streams (Stiber et al., 2022, 2025; Bremers et al., 2023; Spitale et al., 2023). To further explore this potential and benchmark multimodal robot failure detection in HRI, the ERR@HRI 2.0 challenge focuses on the multimodal detection of robot errors and failures in dyadic human-robot conversations.
@@
The ERR@HRI initiative serves as a platform for benchmarking not only HRI datasets but also machine learning models aimed at detecting robot errors and failures from multimedia. The 2024 ERR@HRI challenge, held as part of the 2024 ACM International Conference on Multimodal Interaction (ICMI’24), focused on modeling robot errors using a multimodal dataset collected in a real-world setting where a robotic coach delivered well-being coaching practices to each individual separately over four weeks (Spitale et al., 2024). Building on this, the ERR@HRI 2.0 challenge uses a new multimodal dataset (see Fig. 1) comprised of multimodal behavioral signals extracted from audio and video recordings of human conversations with two distinctive LLM-powered robots: a social robot with more anthropomorphic features and expressions, and a smart speaker. The user and the robot collaborated on five tasks during the conversations, ranging from medical self-diagnosis, trip planning, selecting items more beneficial during a desert survival simulation, to discussing whether the federal government should ban capital punishment and whether universities should have their own police force. Additionally, the ERR@HRI 2.0 challenge is also unique in that it contains two sub-challenges to tackle robot error detection from two perspectives. Sub-challenge 1 targets the detection robot errors and failures from the system perspective, while sub-challenge 2 targets the detection of robot errors and failures from the user perspective.
@@
Through this challenge, we hope to bring together researchers and practitioners from the multimedia community and the robotics and human-robot interaction community, while highlighting the underlying multimedia aspects of the problem. We invite teams to develop and benchmark multimodal machine learning algorithms to detect robot errors and failures in a human-robot conversation dataset, with multimodal features and ground truth labels gathered from approximately 16 hours of multimedia. The techniques developed through this challenge will contribute to advancing our understanding of how multimedia can enhance the analysis and interpretation of interactions between humans and autonomous robots.
@@ 2.1. Robot Errors
As robots become more capable, they are increasingly being integrated into daily life, assisting and collaborating with human users across a range of tasks and contexts (Mahmood et al., 2025a; Gerling et al., 2016; Portugal et al., 2019; Scassellati et al., 2018). However, errors and failures remain common and often inevitable in real-world deployed robots (Carlson and Murphy, 2005; Honig and Oron-Gilad, 2018; Mirnig et al., 2015). For a robot to be truly effective in a human-robot team, it must not only complete its assigned tasks but also adapt to the user and align its behavior with the user’s expectations—thereby enhancing team performance and user satisfaction (Tabrez et al., 2020). This highlights a limitation in existing robot evaluation approaches that only assess the robot based on task completion. Instead, we need to consider how the task was performed and whether the robot’s behavior conformed to their human teammate’s expectations. Therefore, in this work, we adopt a broader definition of robot errors that assesses the robots behavior from two perspective: 1) robot errors from the system perspective—deviations in the robot behavior from its expected behavior by system design; 2) robot errors from the user perspective—deviations in the robot behavior from the user’s mental model of the expected robot behavior.
@@ 2.2. Robot Error Detection Using Social Signals
Social signals are multimodal behavioral cues (both verbal and non-verbal) that can convey information about one’s emotions, intentions, attitudes, and social dynamics within interactions (Vinciarelli et al., 2008; Cao and Huang, 2022; Candon et al., 2023; Sauppé and Mutlu, 2014; Stiber et al., 2024). In human-robot interaction (HRI), users’ natural responses (social signals) during the interaction can impart information about their own internal state (e.g., user uncertainty) and their mental model of the robot to the robot (Vinciarelli et al., 2011; Tabrez et al., 2020; Stiber et al., 2022; Scherf et al., 2024; Huang and Mutlu, 2013). As such, social signals exhibited by users have been used to aid collaboration in HRI by conveying user preferences (e.g., (Candon et al., 2023)), user’s need for help (e.g., (Wilson et al., 2022)), and engagement breakdown (e.g., (Ben-Youssef et al., 2019)).
Social signals have also been used to facilitate automatic error detection in HRI. Robot errors naturally elicit responses from users (Mirnig et al., 2015). Capitalizing on those elicitations expressed as multimodal streams, recent works have shown the possibility using them for detecting robot errors (Stiber et al., 2022, 2023, 2025; Bremers et al., 2023; Spitale et al., 2023; Kontogiorgos et al., 2021; Giuliani et al., 2015; Spitale et al., 2024). The 2024 ERR@HRI challenge (Spitale et al., 2024) at the 2024 ACM International Conference on Multimodal Interaction focused on using a multimodal dataset collected in a real-world setting where a robotic coach delivered well-being coaching practices to each individual separately over four weeks. In an effort to further explore the potential of this and to benchmark multimodal robot failure detection in HRI, the ERR@HRI 2.0 challenge focuses on the multimodal detection of robot errors and failures in dyadic human-robot conversations across two unique embodiments and five task settings.
@@
In continuation of previous year’s ERR@HRI (Spitale et al., 2024) efforts to encourage researchers to develop and benchmark multimodal machine learning models to detect robot errors during human-robot interaction, the ERR@HRI 2.0 challenge invites participants to develop multimodal ML models designed to detect conversational robot failures using a dataset that includes facial, head pose, and speech features extracted from 16 hours of dyadic human-robot conversations multimedia.
3.1. Defining Robot Errors
In this work, we are interested in detecting robot errors both from the system perspective and the user perspective (Section 2). These two types of robot errors overlap but are distinctive as deviations from the system’s designed behavior may not always be errors from the user’s perspective. Similarly, system’s designed behavior that deviates from user’s expected system behaviors are not always errors from the system’s perspective but are considered errors or failures from the user’s perspective. Therefore, it is important to consider both user perceived errors and system perceived errors when creating models that detects errors during human-robot interaction. We operationalize robot errors from the system perspective as a noticable misalignment between robot behavior and its expected behavior by system design. We operationalized robot errors from the user perspective as noticable verbal and non-verbal user-initiated disruptive interruptions that are intended to disrupt and correct for a misalignment between the robot behavior and the user’s expected robot behavior. We want to emphasize that our operationalized definition of robot error used in this work does not capture all robot errors from the system perspective nor the user perspective. However, this definition allows for robot errors to be labeled from a third-party perspective, without having the participants self-report when they felt like the robot made a mistake. Similar approach to how prior work has defined other user internal states, e.g., confusion (Stiber et al., 2024).
3.2. Task
@@
Detection of robot error from the system’s perspective
Detection of robot error from the user’s perspective
@@ 3.3. Dataset
The dataset consists of data from 42 users (21 female, 21 male), in a total of 101 sessions and roughly 950 minutes (15 hours and 50 minutes 48 seconds) of interaction collected in previous dyadic human-social robot conversation study (Cao et al., 2025) and dyadic human-voice assistant conversation study (Mahmood et al., 2025b). Table 1 shows additional characteristics of the dataset.
The human-social robot conversation data consists of 3 hours and 43 minutes 50 seconds of audio and video data of 22 unique participants interacting with a social robot to complete a set of one to two tasks including selected items to aid survival in a desert survival simulation (1 hour 43 minutes 11 seconds) and discussing whether the federal government should ban capital punishment (2 hours 39 seconds). For more details on the task and the study condition under which the data was collected, please refer to the paper (Cao et al., 2025).
The human-voice assistant conversation data consists of 12 hours and 6 minutes 58 seconds of audio and video data, collected from 20 unique participants interacting with the voice assistant on a set of one to three tasks including medical self-diagnosis (3 hours 50 minutes 29 seconds), planning a day trip in Edinburgh (4 hours 40 minutes 45 seconds), and discussing whether universities should have their own police forces (3 hours 35 minutes 44 seconds). For more details on the task and the study condition under which the data was collected, please refer to the paper (Mahmood et al., 2025b).
@@ 3.3.1. Feature Extraction
Facial and head pose features: We used the OpenFace 2.2.0 toolkit (Baltrusaitis et al., 2018) to recognize facial expressions through detecting the presence (18 AUs) and intensity (17 AUs) of facial action unit (AU). We also use the OpenFace 2.2.0 toolkit to estimate the location (x, y, and z coordinates) and rotation (x, y, and z coordinates) of the user’s head. These features combined with the toolkit output on its confidence and success results in a total of 43 features per frame at a rate of 30 frames per second.
@@
Transcribed speech features: We used Google’s speaker diarization to detect user and robot speaker turns in audio recordings. Based on the result, we extracted 515 features showing when the robot/user is speaking, the word count of their speech content, the confidence of the speech-to-text, and created a 512-dimension embedding of the transcribed speech using a ViT-B/16 Transformer architecture Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021) embedding.
@@ 3.4. Metrics
We define an error as detected when the window in which the error is predicted overlaps with the ground truth error duration labeled. Following this definition, a prediction is a true positive when it overlaps with the ground truth error/reactions windows. We added a tolerance of one second to the start and end labels to account for label inaccuracies. If the two windows overlap by any amount then that is considered a true positive. Following this logic, we consider a prediction to be a false positive, when the prediction window does not overlap with any ground truth windows. We used these definitions to evaluate overall model performance.
@@ 3.7. Participation and Conclusion
This paper introduces the ERR@HRI 2.0 Challenge organized in conjunction with the ACM International Conference on Multimedia 2025 (ACM-MM’25), which focuses on detecting robot errors and failures in human-robot conversations. A total of nine teams registered to participate in the challenge. Two teams submitted models to each of the sub-challenges and surpassed the baseline model performance. The two teams will be invited to submit a workshop-style paper describing their ML solutions and results on the dataset, as well as a publicly available code repository. In the future, we hope to continue host more ERR@HRI Grand Challenges in the coming years utilizing diverse human-robot interaction multimedia datasets. Future challenges should explore multimodal detection of different types of errors (physical, cognitive), severity of errors, types of robots (physical manipulator, mobile robots) in different settings (in-the-wild) to further advance multimodal detection of robot errors and failures.
@@ Acknowledgements.
S. Cao, M. Stiber, A. Mahmood & C.-M. Huang’s work was supported by the National Science Foundation award #2141335 and National Science Foundation award #2143704. M. Spitale has been supported by PNRR-PE-AI FAIR project funded by the NextGeneration EU program. H. Gunes has been supported by the EPSRC project ARoEQ under grant ref. EP/R030782/1. This paper has been proofread by a language model (AI), and the authors have read through the resulting content to ensure it accurately reflects the original intent. We would like to acknowledge Yifan Xu for their assistance with the baseline model training and Nithish Krishna Shreenevasan for their contribution to the data labeling.References

### v2

@@ 1. Introduction
In recent years, the integration of large language models (LLMs) into conversational robots—embodied media—has enabled more natural human-robot conversations. While LLMs can mitigate certain speech recognition errors that often cause traditional conversational robots to fail, these advanced systems remain vulnerable to issues such as interrupting users and misinterpreting user intent (Mahmood et al., 2025b, 2024; Cao et al., 2025). Such failures can lead to conversational breakdowns, disrupt interaction flow, and negatively affect users’ perception of the robot. As LLM-powered conversational robots are deployed in real-world scenarios, the need to understand and mitigate their failures is a pressing issue. Prior work has discovered some strategies to help robots rebuild user trust and recover from errors (Karli et al., 2023; Esterwood and Robert, 2021; Axelsson et al., 2024); however, the application of these strategies hinges on the robot’s ability to detect its own errors effectively (Stiber et al., 2023). Status quo robot error detection methods often either depend on task-specific or domain-specific information—which can be hard to generalize across people, tasks, and errors—or rely on users to detect and report errors—which can be delayed. Recent works showed the possibility of detecting robot error through leveraging social signals exhibited by user reactions and responses to robot errors in multimedia streams (Stiber et al., 2025; Bremers et al., 2023; Spitale et al., 2023; Pramanick and Rossi, 2024; Janssens et al., 2024; Wachowiak et al., 2024). To further explore this potential and benchmark multimodal robot failure detection in HRI, the ERR@HRI 2.0 challenge focuses on multimodal detection of robot errors and failures in dyadic human-robot conversations.
@@
The ERR@HRI initiative serves as a platform for benchmarking not only HRI datasets but also machine learning models aimed at detecting robot errors and failures from multimedia. The 2024 ERR@HRI challenge, held as part of the 2024 ACM International Conference on Multimodal Interaction (ICMI’24), focused on modeling robot errors using a multimodal dataset collected in a real-world setting where a robotic coach delivered well-being coaching practices to each individual separately over four weeks (Spitale et al., 2024). Building on this, the ERR@HRI 2.0 challenge uses a new multimodal dataset (see Fig. 1) comprised of multimodal behavioral signals extracted from audio and video recordings of human conversations with two distinctive LLM-powered embodied media: a social robot with more anthropomorphic features and expressions, and a smart speaker. The user and the robot collaborated on five tasks during the conversations, ranging from medical self-diagnosis, trip planning, selecting items more beneficial during a desert survival simulation, to discussing whether the federal government should ban capital punishment and whether universities should have their own police force. Additionally, the ERR@HRI 2.0 challenge is also unique in that it contains two sub-challenges to tackle robot error detection from two perspectives. Sub-challenge 1 targets the detection of robot errors and failures from the system perspective, while sub-challenge 2 targets the detection of robot errors and failures from the user perspective.
@@
Through this challenge, we hope to bring together researchers and practitioners from the multimedia, robotics, and human-robot interaction communities, while highlighting the underlying multimedia aspects of the problem. The techniques developed through this challenge will contribute to advancing our understanding of how multimedia can enhance the analysis and interpretation of interactions between humans and autonomous robots.
@@ 2.1. Robot Errors
As robots become more capable, they are increasingly being integrated into daily life, assisting and collaborating with human users across a range of tasks and contexts (Mahmood et al., 2025a; Gerling et al., 2016; Portugal et al., 2019; Scassellati et al., 2018). However, errors and failures remain common and often inevitable in robots deployed in real-world settings (Carlson and Murphy, 2005; Honig and Oron-Gilad, 2018; Mirnig et al., 2015). For a robot to be truly effective in a human-robot team, it must not only complete its assigned tasks but also adapt to the user and align its behavior with the user’s expectations—thereby enhancing team performance and user satisfaction (Tabrez et al., 2020). This highlights a limitation in existing robot evaluation approaches that only assess the robot based on task completion. Instead, we need to consider how the task was performed and whether the robot’s behavior conformed to their human teammate’s expectations. Therefore, in this work, we adopt a broader definition of robot errors to assess the robot behaviors from two distinct perspectives: 1) robot errors from the system perspective—deviations in the robot behavior from its expected behavior by system design; and 2) robot errors from the user perspective—deviations in the robot behavior from the user’s mental model of the expected robot behavior.
@@ 2.2. Robot Error Detection Using Social Signals
Social signals are multimodal behavioral cues (both verbal and non-verbal) that can convey information about one’s emotions, intentions, attitudes, and social dynamics within interactions (Vinciarelli et al., 2008; Cao and Huang, 2022; Candon et al., 2023; Sauppé and Mutlu, 2014; Stiber et al., 2024). In human-robot interaction (HRI), users’ natural responses (social signals) during the interaction can impart information about their own internal state (e.g., user uncertainty) and their mental model of the robot (Vinciarelli et al., 2011; Tabrez et al., 2020; Stiber et al., 2022; Scherf et al., 2024; Huang and Mutlu, 2013). As such, social signals exhibited by users have been used to aid collaboration in HRI by conveying user preferences (e.g., (Candon et al., 2023)), user’s need for help (e.g., (Wilson et al., 2022)), and engagement breakdowns (e.g., (Ben-Youssef et al., 2019)).
Social signals have also been used to facilitate automatic error detection in HRI. Robot errors naturally elicit responses from users (Mirnig et al., 2015). Capitalizing on those elicitations expressed as multimodal streams, recent works have shown the possibility of using them for detecting robot errors (Stiber et al., 2022, 2023, 2025; Bremers et al., 2023; Spitale et al., 2023; Kontogiorgos et al., 2021; Giuliani et al., 2015; Spitale et al., 2024). The 2024 ERR@HRI challenge (Spitale et al., 2024) at the 2024 ACM International Conference on Multimodal Interaction focused on using a multimodal dataset collected in a real-world setting where a robotic coach delivered well-being coaching practices to each individual separately over four weeks. In an effort to further explore the potential of this and to benchmark multimodal robot failure detection in HRI, the ERR@HRI 2.0 challenge focuses on the multimodal detection of robot errors and failures in dyadic human-robot conversations across two unique embodiments and five tasks.
@@
In continuation of previous year’s ERR@HRI (Spitale et al., 2024) efforts to encourage researchers to develop and benchmark multimodal machine learning models to detect robot errors during human-robot interaction, the ERR@HRI 2.0 challenge invites participants to develop multimodal ML models designed to detect conversational robot failures using a dataset that includes facial, head pose, and speech features extracted from 16 hours of dyadic human-robot conversational media.
3.1. Definition of Robot Errors
In this work, we are interested in detecting robot errors both from the system perspective and the user perspective (Section 2). These two types of robot errors overlap but are distinctive as deviations from the system’s designed behavior may not always constitute errors from the user’s perspective. Similarly, system’s designed behavior that deviates from user’s expected system behaviors are not always errors from the system’s perspective but are considered errors or failures from the user’s perspective. Therefore, it is important to consider both user perceived errors and system perceived errors when creating models that detect errors during human-robot interaction. We operationalize robot errors from the system perspective as a noticable misalignment between robot behavior and its expected behavior by system design. We operationalized robot errors from the user perspective as noticable verbal and non-verbal user-initiated disruptive interruptions that are intended to disrupt and correct for a misalignment between the robot behavior and the user’s expected robot behavior. We want to emphasize that our operationalized definition of robot error used in this work does not capture all robot errors from the system perspective nor the user perspective. However, this definition allows for robot errors to be labeled from a third-party perspective, without having the participants self-report when they felt like the robot made a mistake. A similar approach was used in prior work to define other user internal states, e.g., confusion (Stiber et al., 2024).
3.2. The Tasks
@@
Sub-challenge 1. Detection of robot errors from the system’s perspective
Sub-challenge 2. Detection of robot errors from the user’s perspective
@@ 3.3. Dataset
The dataset consists of data from 42 users (21 female, 21 male), in a total of 101 sessions and roughly 950 minutes (15 hours and 50 minutes 48 seconds) of interaction collected in a previous dyadic human-social robot conversation study (Cao et al., 2025) and dyadic human-voice assistant conversation study (Mahmood et al., 2025b). Table 1 summarizes additional characteristics of the dataset.
The human-social robot conversational data consists of 3 hours and 43 minutes 50 seconds of audio and video data of 22 unique participants interacting with a social robot to complete a set of one to two tasks including selected items to aid survival in a desert survival simulation (1 hour 43 minutes 11 seconds) and discussing whether the federal government should ban capital punishment (2 hours 39 seconds). For more details on the task and the study condition under which the data was collected, please refer to (Cao et al., 2025).
The human-voice assistant conversational data consists of 12 hours and 6 minutes 58 seconds of audio and video data, collected from 20 unique participants interacting with the voice assistant on a set of one to three tasks including medical self-diagnosis (3 hours 50 minutes 29 seconds), planning a day trip in Edinburgh (4 hours 40 minutes 45 seconds), and discussing whether universities should have their own police forces (3 hours 35 minutes 44 seconds). For more details on the task and the study condition under which the data was collected, please refer to the paper (Mahmood et al., 2025b).
@@ 3.3.1. Feature Extraction
Facial and head pose features: We used the OpenFace 2.2.0 toolkit (Baltrusaitis et al., 2018) to identify facial expressions by detecting the presence of 18 facial action units (AUs) and estimating the intensity levels for 17 of those AUs. We also use the OpenFace 2.2.0 toolkit to estimate the location (x, y, and z coordinates) and rotation (x, y, and z coordinates) of the user’s head. These features combined with the toolkit output on its confidence and success results in a total of 43 features per frame at a rate of 30 frames per second.
@@
Transcribed speech features: We used Google’s speaker diarization to detect user and robot speaker turns in audio recordings. Based on the result, we extracted 515 features showing when the robot/user is speaking, the word count of their speech content, the speech-to-text confidence, and created a 512-dimension embedding of the transcribed speech using a ViT-B/16 Transformer architecture Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021) embedding.
@@ 3.4. Metrics
We define an error as detected when the window in which the error is predicted overlaps with the ground truth error duration labeled. Following this definition, a prediction is a true positive when it overlaps with the ground truth error/reaction windows. We added a tolerance of one second to the start and end labels to account for label inaccuracies. If the two windows overlap by any amount then that is considered a true positive. Following this logic, we consider a prediction to be a false positive, when the prediction window does not overlap with any ground truth windows. We used these definitions to evaluate overall model performance.
@@ 3.7. Participation and Conclusion
This paper introduces the ERR@HRI 2.0 Challenge organized in conjunction with the ACM International Conference on Multimedia 2025 (ACM-MM’25), which focuses on detecting robot errors and failures in human-robot conversations. A total of nine teams registered to participate in the challenge. Two teams submitted models to each of the sub-challenges and surpassed the baseline model performance. The two teams will be invited to submit a workshop-style paper describing their ML solutions and results on the dataset, as well as a publicly available code repository. We hope to continue hosting further ERR@HRI Grand Challenges in the coming years utilizing diverse human-robot interaction multimedia datasets. Future challenges should explore multimodal detection of different types of errors (physical, cognitive), severity of errors, types of robots (physical manipulator, mobile robots) in different settings (in-the-wild) to further advance multimodal detection of robot errors and failures.
@@ Acknowledgements.
S. Cao, M. Stiber, A. Mahmood & C. M. Huang’s work was supported by the National Science Foundation award #2141335 and National Science Foundation award #2143704. M. Spitale has been supported by PNRR-PE-AI FAIR project funded by the NextGeneration EU program. H. Gunes has been supported by the EPSRC project ARoEQ under grant ref. EP/R030782/1. This paper has been proofread by a language model (AI), and the authors have read through the resulting content to ensure it accurately reflects the original intent. We would like to acknowledge Yifan Xu for their assistance with the baseline model training and Nithish Krishna Shreenevasan for their contribution to the data labeling.References
@@
- Axelsson et al. (2024) Minja Axelsson, Micol Spitale, and Hatice Gunes. 2024. ”Oh, Sorry, I Think I Interrupted You”: Designing Repair Strategies for Robotic Longitudinal Well-being Coaching. In Proceedings of the 2024 ACM/IEEE International Conference on Human-Robot Interaction. 13–22.
@@
- Brodersen et al. (2010) Kay Henning Brodersen, Cheng Soon Ong, Klaas Enno Stephan, and Joachim M Buhmann. 2010. The balanced accuracy and its posterior distribution. In 2010 20th international conference on pattern recognition. IEEE, 3121–3124.
@@
- Janssens et al. (2024) Ruben Janssens, Eva Verhelst, and Mathieu De Coster. 2024. Predicting Errors and Failures in Human-Robot Interaction from Multi-Modal Temporal Data. In Proceedings of the 26th International Conference on Multimodal Interaction. 671–676.
@@
- Pramanick and Rossi (2024) Pradip Pramanick and Silvia Rossi. 2024. PRISCA at ERR@ HRI 2024: Multimodal Representation Learning for Detecting Interaction Ruptures in HRI. In Proceedings of the 26th International Conference on Multimodal Interaction. 666–670.
@@
- Wachowiak et al. (2024) Lennart Wachowiak, Peter Tisnikar, Andrew Coles, Gerard Canal, and Oya Celiktutan. 2024. A Time Series Classification Pipeline for Detecting Interaction Ruptures in HRI Based on User Reactions. In Proceedings of the 26th International Conference on Multimodal Interaction. 657–665.